OpenStudy (anonymous):

State whether the given measurements determine zero, one, or two triangles. B = 86°, b = 21, c = 18

OpenStudy (rajee_sam):

 Now using rule of sines for a triangle I write\[\frac{ Sin B }{ b } = \frac{ SinC }{ c }\]

OpenStudy (rajee_sam):

\[\frac{ Sin 86 }{ 21 } = \frac{ SinC }{ 18 }\]\[\frac{ .9976 }{ 21 } = \frac{ Sin C }{ 18 }\] Solving this would give you \[Sin C = .885\] \[Angle \space C = 58.77° \] Now sum of angles in a Triangle is 180°. So Angle A = 180-86-58.77 = 35.23 So you can find the third side. This will make only one Triangle

OpenStudy (anonymous):

Alright. I'll remember this. Thank you.